# roundedstyle backgroundcolor=verylightgray, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt, linewidth=1pt

### v2

@@ Abstract
Self-correction in text-to-SQL is the process of prompting large language model (LLM) to revise its previously incorrectly generated SQL, and commonly relies on manually crafted self-correction guidelines by human experts that are not only labor-intensive to produce but also limited by the human ability in identifying all potential error patterns in LLM responses. We introduce MAGIC, a novel multi-agent method that automates the creation of the self-correction guideline. MAGIC uses three specialized agents: a manager, a correction, and a feedback agent. These agents collaborate on the failures of an LLM-based method on the training set to iteratively generate and refine a self-correction guideline tailored to LLM mistakes, mirroring human processes but without human involvement. Our extensive experiments show that MAGIC’s guideline outperforms expert human’s created ones. We empirically find out that the guideline produced by MAGIC enhance the interpretability of the corrections made, providing insights in analyzing the reason behind the failures and successes of LLMs in self-correction. We make all agent interactions publicly available to the research community, to foster further research in this area, offering a synthetic dataset for future explorations into automatic self-correction guideline generation.
1 Introduction
Converting natural language questions to SQL database queries, known as text-to-SQL, serves as a pivotal component for empowering non-expert data analysts in extracting desired information from relational databases using natural language Qu et al. (2024). While large language models have shown a significant improvement in text-to-SQL and serve as state-of-the methods according to the leaderboards Li et al. (2024b), they are prone to mistakes and even GPT4 has a notable accuracy gap of 30% within human Li et al. (2023).
A solution for resolving the mistakes of LLM is the emerging concept of ‘self-correction’ Madaan et al. (2023) that defines as the ability of LLMs in revising their previous mistakes under the hypothesis that recognizing errors is easier than avoiding them Gou et al. (2024); Madaan et al. (2023). Self-correction in context of text-to-SQL is the process of prompting an LLM to revise its previously incorrectly generated SQL Zhang et al. (2024); Chen et al. (2023).
@@ 1 Introduction
While self-correction for LLMs has been widely studied in variety of tasks including code generation with GPT4 Shypula et al. (2024); Pan et al. (2023); Kamoi et al. (2024), it has been relatively unexplored in text-to-SQL. The common method for self-correcting in existing few-shot LLM based methods that achieve state-of-the-art effectiveness is designing a self-correction guideline by human and prompting it to GPT-4 together with its initial generated SQL Pourreza et al. (2023). This self-correction guideline commonly engineered based on the train set where an expert human analyzes common mistakes of the LLM on the training data and design a guideline to prevent common mistakes. This process is a time-consuming and challenging task that is limited to the ability of humans in identifying all the mistake patterns that exist in the LLM responses Pourreza et al. (2023); Wang et al. (2024a); Talaei et al. (2024).
@@
Unlike previous studies that analyze self-correction on a simple method that uses GPT4 as its backbone or a low-effective open-source language model Zhang et al. (2024), we employ a strong and effective open-source method based on GPT4, DIN-SQL Pourreza et al. (2023), as our initial text-to-SQL system aiming to compare our automatically generated guideline with an effective expert human-written guideline. Our experiments demonstrate that the generated guideline by MAGIC leads to self-explanatory self-correction, where the LLM asks itself questions before self-correcting. Figure 1 illustrates an example of where the MAGIC guideline assisted GPT4 to self-correct its previously incorrectly generated SQL. The process is interpretable as the LLM first starts with asking itself and then perform self-correction.
@@
  • We establish the task of self-correction guideline generation for text-to-SQL, and introduce MAGIC, a novel multi-agent method, that automates the generation of self-correction guideline and outperforms human-written guideline; yielding to improve effectiveness of a strong few-shot LLM based text-to-SQL method.
  • We systematically analyze the impact of ours and existing self-correction methods to perform self-correction across different scenarios: Correcting incorrect queries; non-executable queries; and all queries.
  • We reproduce all the baselines utilizing the same version of GPT-4 across all the experiments, and provide a comprehensive comparable insight on self-correction in text-to-SQL.
@@
We found that there can be a significant gap in terms of effectiveness when utilizing different versions of GPT-4. For instance, our replication of DIN-SQL Pourreza et al. (2023) on development set of BIRD Li et al. (2023) achieves 56.52 compared to the original report of 50.72 in the paper, which is about 6 points higher than the reported number. This is our motivation for replicating and reproducing all the reported baselines using same GPT-4 and and avoid copying numbers from previous papers.
2 Related work
@@
LLMs for Text-to-SQL. Converting natural language questions into SQL queries, known as text-to-SQL, has been an active area of research within both the natural language processing (NLP) and database communities for many years Zelle and Mooney (1996); Guo et al. (2019). Recently, text-to-SQL has benefited from the promising effectiveness of Large Language Models (LLMs) Talaei et al. (2024); Pourreza et al. (2023). Early methods utilized the zero-shot in-context learning capabilities of LLMs for SQL generation Rajkumar et al. (2022). Building on this, subsequent models have improved LLM performance through task decomposition and techniques such as Chain-of-Thought (CoT) Wei et al. (2022), including but not limited to models like DAIL-SQL Gao et al. (2023), MAC-SQL Wang et al. (2024a), C3 Dong et al. (2023), self-consistency Wang et al. (2022), and least-to-most prompting Zhou et al. (2023). In addition, there are studies focusing on fine-tuning LLMs for text-to-SQL Li et al. (2024c); Gao et al. (2023); Li et al. (2024a). We choose DIN-SQL Pourreza et al. (2023) as the initial text-to-SQL system due to two reasons: (i) high effectiveness and available open-source code; (ii) its self-correction component is a human expert written guideline based on prior mistakes of the LLM making it a suitable baseline for comparing the automatically self-correction guideline generated by MAGIC with expert human-written guideline. However, we would emphasize that our method is independent of how the in-context learning method is designed and is adaptable to the output of any method, as we only require the predicted SQL of the initial text-to-SQL system for tackling self-correction in MAGIC.
@@
Self-correction in text-to-SQL. Pan et al. (2024) and Kamoi et al. (2024) provide an extensive overview of research on self-correction in a variety of domains up to 2024. Here, we focus on self-correction in Text-to-SQL, where there has been relatively limited study. Self-debugging Chen et al. (2023) generates additional explanations on the question and initial predicted SQL, which are then provided to an LLM for self-correction. DIN-SQL Pourreza et al. (2023) designs a human-written self-correction guideline and revises all the initially generated SQLs according to this guideline. Self-consistency is based on generating several candidates and employing a voting process, which has been integrated into Text-to-SQL by DAIL-SQL Gao et al. (2023). The refiner component of MAC-SQL Wang et al. (2024a) uses execution errors as signals for revising the SQL. To the best of our knowledge, we are the first study to address self-correction in Text-to-SQL by generating a self-correction guideline.
@@
LLM-based Agents. LLM-based agents have been a promising area of study in both academic and industry communities for an extended period Wang et al. (2024b), leading to extensive research exploring autonomous agents based on LLMs such as AutoGPT Significant Gravitas , OpenAgents Xie et al. (2023), and AutoGen Wu et al. (2023). However, there are limited studies leveraging this concept for text-to-SQL, with MAC-SQL being the only multi-agent method focusing on addressing the Text-to-SQL task through a new multi-agent collaborative framework Wang et al. (2024a). Inspired by the literature on multi-agent LLMs, in MAGIC, we employ a multi-agent collaborative framework that iteratively analyzes the failures of predicted SQLs by a text-to-SQL method’s and automatically generates a self-correction guideline tailored to the method’s mistakes. To the best of our knowledge, there is no prior work in the literature that employs multi-agent methods for designing a self-correction guideline in text-to-SQL or other domains.
3 Self-correction Guideline Generation
@@
Tasks definition. Given a set of questions in natural language alongside the corresponding database schema and the initially generated incorrect SQL queries by a text-to-SQL system, the task of self-correction guideline generation is to generate a self-correction guideline that is tailored to the LLM’s mistakes and prevents it from repeating them. This can be done by a human expert, an LLM, or a collaboration between a human expert and an LLM. The typical approach for self-correction involves writing a guideline for the Large Language Model manually by humans based on the LLM’s common mistakes Pourreza et al. (2023); Talaei et al. (2024). The data used for guideline generation must not be shared with the evaluation data to prevent overfitting the guidelines on the test errors.
3.1 MAGIC
@@
Our method, MAGIC, consists of three agents illustrated in Figure 2: the manager, feedback, and correction agents. These agents interact iteratively over the failures to generate the self-correction guideline. We describe this process in detail in the following.
@@
Revising agents’ instruction. In the first iteration of the feedback-correction cycle, the manager uses two predefined prompts designed for its interactions with the feedback and correction agents, as illustrated in Figures 5 and 6 respectively. If the correction agent cannot successfully revise $s^{\prime}$ in the first iteration, the manager begins revising these predefined prompts. The prompt template of manager for revising the predefined prompts of feedback and correction agents are illustrated in Figures 7 and 8. We found empirically that this step plays an important role for the manager to adapt its interaction with agents. Manager revise the agent’s instruction based on the previous response of the agent, previous prompt that is used for interacting with the agent, question, $s^{\prime}$, and $s^{gt}$.
Guideline generation. Given each successful revision by the correction agent, the manager stores the corresponding successful feedback that led to this in its memory. The manager then aggregates these stored feedbacks to generate the self-correction guideline batch-by-batch using a predefined prompt for guideline generation (Figure 9). Each batch consist of $k$ feedbacks. In our preliminary experiments, we determined that a feedback batch size of 10 is optimal. In the first batch, the manager begins without any pre-existing guideline and generate an initial guideline. For subsequent batches, the manager updates the already generated guideline. This guideline generation process is triggered at the end of a feedback-correction cycle if the manager has accumulated a new batch of successful feedbacks that has not yet been utilized for guideline generation. The self-correction guideline generated by our proposed method, MAGIC, automatically and from scratch, is presented in Figures 11, 12, 13, 14, 15, and 16 in appendix section.
Efficiency. Previous approaches to craft self-correction guidelines needed extensive expert human work. With MAGIC, we can efficiently generate, empirically shown better guideline, in less than 2 hour.
3.2 Agents access to information
@@
Self-Debugging Chen et al. (2023) 63.57 50.11 39.67 57.24 63.35 49.03 39.58 56.78 63.03 49.25 38.89 56.58
Self-Consistency Wang et al. (2022) 64.56 50.79 40.09 58.08 63.24 49.03 38.19 56.58 64.00 49.68 37.50 57.17
Multiple-Prompt Lee et al. (2024) 64.22 50.75 40.02 57.86 63.35 49.03 38.19 56.65 63.68 50.11 39.58 57.30
@@
Human Expert G Wang et al. (2024a) 63.18 49.19 38.30 56.60 63.20 49.03 38.19 56.58 47.80 46.22 35.30 46.14
Human Expert G Pourreza et al. (2023) 64.32 50.32 39.58 57.76 63.58 49.03 38.89 56.98 62.49 48.17 37.50 55.80
@@
Table 1: Effectiveness results in terms of Execution Accuracy (EX) on the DEV set of the BIRD dataset. All the experiments have been reproduced by us. ‘S’, ‘M’, ‘C’, and ‘T’ refer to different levels of difficulty: Simple, Medium, Challenging, and Total, respectively. The baseline for Natural Language to SQL conversion is DIN-SQL and ‘Human Expert G’ refers to self-correction guideline (prompt) by Pourreza et al. (2023) and Wang et al. (2024a). ‘MAGIC refers to the guideline that is automatically generated by our proposed method, MAGIC, using the train set of BIRD dataset.
4 Experimental setup
@@
Datasets. The Spider Yu et al. (2018) dataset is a collection of 10,181 questions and 5,693 unique complex SQL queries across 200 databases in 138 domains, with each domain featuring multiple tables. It is divided into training, development, and test sets with 8,659, 1,034, and 2,147 examples, respectively, across 146, 20, and 34 distinct databases, ensuring no overlap between sets. Queries are classified into four difficulty levels based on complexity factors such as SQL keywords, nested subqueries, and use of column selections and aggregations. The BIRD dataset Li et al. (2023) comprises 12,751 unique question-SQL pairings across 95 relative large databases (33.4 GB) in 37 professional domains like blockchain and healthcare. BIRD introduces external knowledge as an additional resource for generating accurate SQL queries to bring more complexity into the task.
@@
Metrics. We employ two key metrics from existing literature: Execution Accuracy (EX) and Valid Efficiency Score (VES). Execution Accuracy measures the correctness of a predicted SQL query by comparing its execution output with that of the ground truth SQL query. A SQL query is deemed correct if its execution results match those of the ground truth, allowing for a precise assessment of the model’s performance given the possibility of multiple valid SQL queries for a single question. The Valid Efficiency Score evaluates the efficiency of executing the generated SQL queries, focusing on both their accuracy and execution time, but only considers queries that produce correct results, i.e., those whose execution outcomes are consistent with the reference query. This dual-metric approach provides a comprehensive evaluation of the model’s ability to generate both accurate and efficient SQL queries.
@@
Baselines. We reproduce DIN-SQL using the publicly available original implementation Pourreza et al. (2023). For self-correction, we employ an extensive set of baselines categorized into guideline-dependent and guideline-independent methods. For guideline-independent methods, we reproduce self-debugging Chen et al. (2023) based on the written prompts in the paper. For self-consistency Wang et al. (2022); Gao et al. (2023), we generate 20 SQL queries given the SQL generation prompt of DIN-SQL instead of generating only one SQL and select the final SQL based on voting. In voting, we execute all SQL queries and select the most frequently returned result’s SQL. If one result is produced by various SQL queries, we select the most efficient one by comparing them against each other. For the Multiple-Prompt baseline, we follow the approach in Lee et al. (2024) by re-ordering candidate tables in the prompt and generating up to 20 different combinations, employing a voting mechanism similar to our self-consistency implementation. For human-expert baselines, we utilize the self-correction guideline from DIN-SQL Pourreza et al. (2023), written by human experts, and the revision guideline from MAC-SQL Wang et al. (2024a), that is also written by human experts and are designed specifically for correcting SQL queries with execution errors. However, we also adopt their approach for correcting incorrect SQL queries and for correcting all initially predicted SQL queries, regardless of whether they are correct, similar to the scenarios analyzed in Chen et al. (2023) and Pourreza et al. (2023) respectively.
5 Results

### v3

@@ Abstract
Self-correction in text-to-SQL is the process of prompting large language model (LLM) to revise its previously incorrectly generated SQL, and commonly relies on manually crafted self-correction guidelines by human experts that are not only labor-intensive to produce but also limited by the human ability in identifying all potential error patterns in LLM responses. We introduce MAGIC, a novel multi-agent method that automates the creation of the self-correction guideline. MAGIC uses three specialized agents: a manager, a correction, and a feedback agent. These agents collaborate on the failures of an LLM-based method on the training set to iteratively generate and refine a self-correction guideline tailored to LLM mistakes, mirroring human processes but without human involvement. Our extensive experiments show that MAGIC’s guideline outperforms expert human’s created ones. We empirically find out that the guideline produced by MAGIC enhances the interpretability of the corrections made, providing insights in analyzing the reason behind the failures and successes of LLMs in self-correction. All agent interactions are publicly available at https://huggingface.co/datasets/microsoft/MAGIC.
Introduction
Converting natural language questions to SQL database queries, known as text-to-SQL, serves as a pivotal component for empowering non-expert data analysts in extracting desired information from relational databases using natural language (Qu et al. 2024). While large language models have shown a significant improvement in text-to-SQL and serve as state-of-the methods according to the leaderboards (Li et al. 2024c), they are prone to mistakes and even GPT4 has a notable accuracy gap of 30% within human (Li et al. 2023).
A solution for resolving the mistakes of LLM is the emerging concept of ‘self-correction’ (Madaan et al. 2023) that defines as the ability of LLMs in revising their previous mistakes under the hypothesis that recognizing errors is easier than avoiding them (Gou et al. 2024; Madaan et al. 2023). Self-correction in context of text-to-SQL is the process of prompting an LLM to revise its previously incorrectly generated SQL (Zhang et al. 2024; Chen et al. 2023).
@@ Introduction
While self-correction for LLMs has been widely studied in variety of tasks including code generation with GPT4 (Shypula et al. 2024; Pan et al. 2023; Kamoi et al. 2024), it has been relatively unexplored in text-to-SQL. The common method for self-correcting in existing few-shot LLM based methods that achieve state-of-the-art effectiveness is designing a self-correction guideline by human and prompting it to GPT-4 together with its initial generated SQL (Pourreza and Rafiei 2023). This self-correction guideline commonly engineered based on the train set where an expert human analyzes common mistakes of the LLM on the training data and design a guideline to prevent common mistakes. This process is a time-consuming and challenging task that is limited to the ability of humans in identifying all the mistake patterns that exist in the LLM responses (Pourreza and Rafiei 2023; Wang et al. 2024a; Talaei et al. 2024).
@@
Unlike previous studies that analyze self-correction on a simple method that uses GPT4 as its backbone or a low-effective open-source language model (Zhang et al. 2024), we employ a strong and effective open-source method based on GPT4, DIN-SQL (Pourreza and Rafiei 2023), as our initial text-to-SQL system aiming to compare our automatically generated guideline with an effective expert human-written guideline. Our experiments demonstrate that the generated guideline by MAGIC leads to self-explanatory self-correction, where the LLM asks itself questions before self-correcting. Figure 1 illustrates an example of where the MAGIC guideline assisted GPT4 to self-correct its previously incorrectly generated SQL. The process is interpretable as the LLM first starts with asking itself and then perform self-correction.
@@
Our contributions are as follows: (i) We establish the task of self-correction guideline generation for text-to-SQL, and introduce MAGIC, a novel multi-agent method, that automates the generation of self-correction guideline and outperforms human-written guideline; yielding to improve effectiveness of a strong few-shot LLM based text-to-SQL method. (ii) We systematically analyze the impact of ours and existing self-correction methods to perform self-correction across different scenarios: Correcting incorrect queries; non-executable queries; and all queries. (iii) We reproduce all the baselines utilizing the same version of GPT-4 across all the experiments, and provide a comprehensive comparable insight on self-correction in text-to-SQL. (iv) We publish all code to reproduce our experiments as open source. [^1] https://github.com/microsoft/SynQo
@@
We found that there can be a significant gap in terms of effectiveness when utilizing different versions of GPT-4. For instance, our replication of DIN-SQL (Pourreza and Rafiei 2023) on development set of BIRD (Li et al. 2023) achieves 56.52 compared to the original report of 50.72 in the paper, which is about 6 points higher than the reported number. This is our motivation for replicating and reproducing all the reported baselines using same GPT-4 and avoid copying numbers from previous papers.
Related work
@@
LLMs for Text-to-SQL. Converting natural language questions into SQL queries, known as text-to-SQL, has been an active area of research within both the natural language processing (NLP) and database communities for many years (Zelle and Mooney 1996; Guo et al. 2019). Recently, text-to-SQL has benefited from the promising effectiveness of Large Language Models (LLMs) (Talaei et al. 2024; Pourreza and Rafiei 2023). Early methods utilized the zero-shot in-context learning capabilities of LLMs for SQL generation (Rajkumar, Li, and Bahdanau 2022). Building on this, subsequent models have improved LLM performance through task decomposition and techniques such as Chain-of-Thought (CoT) (Wei et al. 2022), including but not limited to models like DAIL-SQL (Gao et al. 2023), MAC-SQL (Wang et al. 2024a), C3 (Dong et al. 2023), self-consistency (Wang et al. 2022), and least-to-most prompting (Zhou et al. 2023). In addition, there are studies focusing on fine-tuning LLMs for text-to-SQL (Li et al. 2024d; Gao et al. 2023; Li et al. 2024b). We choose DIN-SQL (Pourreza and Rafiei 2023) as the initial text-to-SQL system due to two reasons: (i) high effectiveness and available open-source code; (ii) its self-correction component is a human expert written guideline based on prior mistakes of the LLM making it a suitable baseline for comparing the automatically self-correction guideline generated by MAGIC with expert human-written guideline. However, we would emphasize that our method is independent of how the in-context learning method is designed and is adaptable to the output of any method, as we only require the predicted SQL of the initial text-to-SQL system for tackling self-correction in MAGIC.
@@
Self-correction in text-to-SQL. Pan et al. (2024) and (Kamoi et al. 2024) provide an extensive overview of research on self-correction in a variety of domains up to 2024. Here, we focus on self-correction in Text-to-SQL, where there has been relatively limited study. Self-debugging (Chen et al. 2023) generates additional explanations on the question and initial predicted SQL, which are then provided to an LLM for self-correction. DIN-SQL (Pourreza and Rafiei 2023) designs a human-written self-correction guideline and revises all the initially generated SQLs according to this guideline. Self-consistency is based on generating several candidates and employing a voting process, which has been integrated into Text-to-SQL by DAIL-SQL (Gao et al. 2023). The refiner component of MAC-SQL (Wang et al. 2024a) uses execution errors as signals for revising the SQL. To the best of our knowledge, we are the first study to address self-correction in Text-to-SQL by generating a self-correction guideline.
@@
LLM-based Agents. LLM-based agents have been a promising area of study in both academic and industry communities for an extended period (Wang et al. 2024b), leading to extensive research exploring autonomous agents based on LLMs such as AutoGPT (Significant Gravitas ), OpenAgents (Xie et al. 2023), and AutoGen (Wu et al. 2023). However, there are limited studies leveraging this concept for text-to-SQL, with MAC-SQL being the only multi-agent method focusing on addressing the Text-to-SQL task through a new multi-agent collaborative framework (Wang et al. 2024a). Inspired by the literature on multi-agent LLMs, in MAGIC, we employ a multi-agent collaborative framework that iteratively analyzes the failures of predicted SQLs by a text-to-SQL method and automatically generates a self-correction guideline tailored to the method’s mistakes. To the best of our knowledge, there is no prior work in the literature that employs multi-agent methods for designing a self-correction guideline in text-to-SQL or other domains.
Self-correction Guideline Generation
@@
Tasks definition. Given a set of questions in natural language alongside the corresponding database schema and the initially generated incorrect SQL queries by a text-to-SQL system, the task of self-correction guideline generation is to generate a self-correction guideline that is tailored to the LLM’s mistakes and prevents it from repeating them. This can be done by a human expert, an LLM, or a collaboration between a human expert and an LLM. The typical approach for self-correction involves writing a guideline for the Large Language Model manually by humans based on the LLM’s common mistakes (Pourreza and Rafiei 2023; Talaei et al. 2024). The data used for guideline generation must not be shared with the evaluation data to prevent overfitting the guidelines on the test errors.
MAGIC
@@
Our method, MAGIC, consists of three agents illustrated in Figure 2: the manager, feedback, and correction agents. These agents interact iteratively over the failures to generate the self-correction guideline. We describe this process in detail in the following. Please note that due to space limitations, the paper includes summarized prompts. We provide the complete prompts and the full automatically generated self-correction guideline by MAGIC in the technical appendix.
@@
Revising agents’ instruction. In the first iteration of the feedback-correction cycle, the manager uses two predefined prompts designed for its interactions with the feedback and correction agents, as illustrated in Figures 3 and 4 respectively. If the correction agent cannot successfully revise $s^{\prime}$ in the first iteration, the manager begins revising these predefined prompts. The prompt template of manager for revising the predefined prompts of feedback and correction agents are illustrated in Figure 5. We found empirically that this step plays an important role for the manager to adapt its interaction with agents. Manager revise the agent’s instruction based on the previous response of the agent, previous prompt that is used for interacting with the agent, question, $s^{\prime}$, and $s^{gt}$.
Guideline generation. Given each successful revision by the correction agent, the manager stores the corresponding successful feedback that led to this in its memory. The manager then aggregates these stored feedbacks to generate the self-correction guideline batch-by-batch using a predefined prompt for guideline generation (Figure 6). Each batch consist of $k$ feedbacks. In our preliminary experiments, we determined that a feedback batch size of 10 is optimal. In the first batch, the manager begins without any pre-existing guideline and generate an initial guideline. For subsequent batches, the manager updates the already generated guideline. This guideline generation process is triggered at the end of a feedback-correction cycle if the manager has accumulated a new batch of successful feedbacks that has not yet been utilized for guideline generation. The self-correction guideline generated by our proposed method, MAGIC, automatically and from scratch, is presented in the technical appendix.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] "question": "{question}",
"Correct SQL": "{Correct SQL}",
"Incorrect SQL": "{Incorrect SQL}". The mistakes are:
Figure 3: The predefined prompt for interaction of manager with feedback agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] - Schema Overview: {schema}
- Question: {question}
- Predicted SQL: “‘sql {Incorrect SQL} “‘
- Expert Human Feedback: {feedback}
Figure 4: The predefined prompt for interaction of manager with correction agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] Manager, review the following prompt for the agent and generate a revised prompt.
Agent description: {Agent description}.
Previous output and prompt that was not useful: {agent_outputs[-1]} and {agent_prompts[-1]}
Revise prompt (Return the entire prompt):
Figure 5: The prompt template of manager for revising the predefined prompts of feedback or correction agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] # Recent mistakes to add to Guideline:
{batch_of_successful_feedbacks}
# Updated Guideline (Return the entire guideline):
Figure 6: The prompt of manager for self-correction guideline generation.
Efficiency. Previous approaches to craft self-correction guidelines needed extensive expert human work. With MAGIC, we can efficiently generate, empirically shown better guideline, in less than 2 hours.
Agents access to information
@@
Self-Debugging (Chen et al. 2023) 63.57 50.11 39.67 57.24 63.35 49.03 39.58 56.78 63.03 49.25 38.89 56.58
Self-Consistency (Wang et al. 2022) 64.56 50.79 40.09 58.08 63.24 49.03 38.19 56.58 64.00 49.68 37.50 57.17
Multiple-Prompt (Lee et al. 2024) 64.22 50.75 40.02 57.86 63.35 49.03 38.19 56.65 63.68 50.11 39.58 57.30
@@
Human Expert G, MAC-SQL (Wang et al. 2024a) 63.18 49.19 38.30 56.60 63.20 49.03 38.19 56.58 47.80 46.22 35.30 46.14
Human Expert G, DIN-SQL (Pourreza and Rafiei 2023) 64.32 50.32 39.58 57.76 63.58 49.03 38.89 56.98 62.49 48.17 37.50 55.80
@@
Table 1: Effectiveness results in terms of Execution Accuracy (EX) on the DEV set of the BIRD dataset. All the experiments have been reproduced by us. ‘S’, ‘M’, ‘C’, and ‘T’ refer to different levels of difficulty: Simple, Medium, Challenging, and Total, respectively. The baseline for Natural Language to SQL conversion is DIN-SQL and ‘Human Expert G’ refers to self-correction guideline (prompt). ‘MAGIC refers to the guideline that is automatically generated by our proposed method, MAGIC, using the train set of BIRD dataset. Significance is shown with $\dagger$ for MAGIC compared to both ‘Human Expert’ G baseliens. Statistical significance was measured with a paired t-test ($p<0.05$) with Bonferroni correction for multiple testing.
Experimental setup
@@
Datasets. The Spider (Yu et al. 2018) dataset is a collection of 10,181 questions and 5,693 unique complex SQL queries across 200 databases in 138 domains, with each domain featuring multiple tables. It is divided into training, development, and test sets with 8,659, 1,034, and 2,147 examples, respectively, across 146, 20, and 34 distinct databases, ensuring no overlap between sets. Queries are classified into four difficulty levels based on complexity factors such as SQL keywords, nested subqueries, and use of column selections and aggregations. The BIRD dataset (Li et al. 2023) comprises 12,751 unique question-SQL pairings across 95 relative large databases (33.4 GB) in 37 professional domains like blockchain and healthcare. BIRD introduces external knowledge as an additional resource for generating accurate SQL queries to bring more complexity into the task.
@@
Metrics. We employ two key metrics from existing literature: Execution Accuracy (EX) and Valid Efficiency Score (VES). Execution Accuracy measures the correctness of a predicted SQL query by comparing its execution output with that of the ground truth SQL query. A SQL query is deemed correct if its execution results match those of the ground truth, allowing for a precise assessment of the model’s performance given the possibility of multiple valid SQL queries for a single question. The Valid Efficiency Score evaluates the efficiency of executing the generated SQL queries, focusing on both their accuracy and execution time, but only considers queries that produce correct results, i.e., those whose execution outcomes are consistent with the reference query. This dual-metric approach provides a comprehensive evaluation of the model’s ability to generate both accurate and efficient SQL queries. Given our primary focus on enhancing accuracy, we prioritize Execution Accuracy (EX) as the main metric in our experiments, while also reporting the Valid Efficiency Score (VES) in selected cases to offer additional insights. Furthermore, we do not report VES results for the SPIDER dataset because computing VES for this dataset is less meaningful. The execution times for SPIDER are so short that any time differences often round to zero, rendering VES comparisons inconclusive. This is why previous studies also omit VES for SPIDER dataset (Pourreza and Rafiei 2023; Talaei et al. 2024). The SPIDER dataset’s databases contain significantly fewer entries compared to the BIRD dataset, where the databases are, on average, 140 times larger. This substantial difference in scale makes VES a more relevant metric for BIRD, where efficiency considerations are more significant.
@@
Baselines. We reproduce DIN-SQL using the publicly available original implementation (Pourreza and Rafiei 2023). For self-correction, we employ an extensive set of baselines categorized into guideline-dependent and guideline-independent methods. For guideline-independent methods, we reproduce self-debugging (Chen et al. 2023) based on the written prompts in the paper. For self-consistency (Wang et al. 2022; Gao et al. 2023), we generate 20 SQL queries given the SQL generation prompt of DIN-SQL instead of generating only one SQL and select the final SQL based on voting. In voting, we execute all SQL queries and select the most frequently returned result’s SQL. If one result is produced by various SQL queries, we select the most efficient one by comparing them against each other. For the Multiple-Prompt baseline, we follow the approach in (Lee et al. 2024) by re-ordering candidate tables in the prompt and generating up to 20 different combinations, employing a voting mechanism similar to our self-consistency implementation. For human-expert baselines, we utilize the self-correction guideline from DIN-SQL (Pourreza and Rafiei 2023), written by human experts, and the revision guideline from MAC-SQL (Wang et al. 2024a), that is also written by human experts and are designed specifically for correcting SQL queries with execution errors. However, we also adopt their approach for correcting incorrect SQL queries and for correcting all initially predicted SQL queries, regardless of whether they are correct, similar to the scenarios analyzed in (Chen et al. 2023) and (Pourreza and Rafiei 2023) respectively. Our method, MAGIC, is designed to generate self-correction guidelines for text-to-SQL tasks without relying on fine-tuned models. We exclude baselines that depend on fine-tuning (Talaei et al. 2024; Pourreza and Rafiei 2024; Li et al. 2024a) to ensure fair comparisons and to highlight MAGIC’s effectiveness in scenarios where fine-tuning is not feasible due to high computational costs.
Results